# Formation of Representations in Neural Networks

### v1

@@ Abstract
Understanding neural representations will help open the black box of neural networks and advance our scientific understanding of modern AI systems. However, how complex, structured, and transferable representations emerge in modern neural networks has remained a mystery. Building on previous results, we propose the Canonical Representation Hypothesis (CRH), which posits a set of six alignment relations to universally govern the formation of representations in most hidden layers of a neural network. Under the CRH, the latent representations (R), weights (W), and neuron gradients (G) become mutually aligned during training. This alignment implies that neural networks naturally learn compact representations, where neurons and weights are invariant to task-irrelevant transformations. We then show that the breaking of CRH leads to the emergence of reciprocal power-law relations between R, W, and G, which we refer to as the Polynomial Alignment Hypothesis (PAH). We present a minimal-assumption theory demonstrating that the balance between gradient noise and regularization is crucial for the emergence the canonical representation. The CRH and PAH lead to an exciting possibility of unifying major key deep learning phenomena, including neural collapse and the neural feature ansatz, in a single framework.
@@ 1 Introduction
The success of deep learning is often attributed to its ability to learn latent representations from data (Bengio et al.,, 2013). These latent representations, progressively formed as data passes through the network’s layers, are found to encode increasingly abstract features of the input:
@@
In this work, we seek to bridge this gap by introducing the Canonical Representation Hypothesis (CRH). At its core, the CRH posits that neural networks, during training, inherently align their representations with the gradients and weights. Furthermore, the satisfaction and breaking of a subset of CRH equations are found to delineate the universal phases, which are empirically observable scaling relationships between the weights, activations, and gradients. The CRH reveals a striking aspect of representation learning: there may exist a set of universal equations that govern the formation of representations and universal phases which distinguish the layers in modern neural networks, independent of the task, architecture, or loss function. Thus, the CRH provides an useful perspective on how neural networks evolve toward compact and interpretable solutions.
@@
    Proposal and justification of the CRH which states that within a layer, the neuron gradients, latent representation, and parameters are driven into mutual alignment after training, due to noise regularization balance (Section 3 and 4);
@@ 2 Related Works
Empirical results show that the representations of well-trained neural networks share universal characteristics (Maheswaranathan et al.,, 2019; Huh et al.,, 2024). For us, a closely related phenomenon is the neural collapse (NC) (Papyan et al.,, 2020), which studies how structured and noise-robust low-rank representations emerge in a classification model. The CRH can be seen as a generalization of NC because one can prove that when restricted to certain settings, the CRH is equivalent to the NC (Section 7). Another related phenomenon is the neural feature ansatz (NFA) (Radhakrishnan et al.,, 2023), which shows that the weight matrices of fully connected layers evolve during training. However, the NFA studies the weight evolution, not the representations. Empirical power-laws are known to exist in large neural networks (Kaplan et al.,, 2020; Bahri et al.,, 2024), which relates the model performance to their sizes. The power laws discovered in our work are different because they are reciprocal relations that relate dual objects (R, G, W) to each other rather than to the performance. Other related works are discussed in the context where they become relevant. Also, due to space constraints, more related works regarding neural representations are discussed in Appendix A.
@@ 3 Canonical Representation Hypothesis
Much recent literature has suggested that the quantities $h_{a},\ h_{b},\ W$ and their gradients are correlated with each other after or throughout training. The neural collapse phenomenon suggests that in a deep overparameterized classifier, $\mathbb{E}[h_{a}h_{a}^{\top}]\propto W^{\top}W$ for the penultimate fully connected layer (Papyan et al.,, 2020; Xu et al., 2023b, ; Ji et al.,, 2021; Kothapalli,, 2022; Rangamani et al.,, 2023). In the study of kernel and feature learning, a primary mechanism of how the neural tangent kernel changes is that after a few steps of update, the representations become correlated with the weights (Everett et al.,, 2024) and so the quantity $Wh_{a}$ will be significantly away from zero, which also implies a strong relationship between $W^{\top}W$ and $\mathbb{E}[h_{a}h_{a}^{\top}]$. The recent work on neural feature ansatz shows that $W^{\top}W\propto\nabla_{h_{a}}f\nabla_{h_{a}}^{\top}f$ for fully connected networks (Radhakrishnan et al.,, 2023; Gan and Poggio,, 2024). The idea that the latent variables will become correlated with the weight updates is also a central notion in the feature learning literature (Yang and Hu,, 2020; Everett et al.,, 2024).
@@
For the formation of representations in neural networks, the most important relations in Eq. (3)-(5) are perhaps the forward and backward RGA, as they directly relate the representations $\mathbb{E}[hh^{\top}]$ to their gradients with respect to the loss function. Because there is no scientific reason for us to believe that the forward representation is more important than the backward representation or vice versa, one should study both directions carefully. The backward RWA in its general form has not been discussed in the literature but has been implicitly studied in the particular setting of neural collapse, which happens in the penultimate layer of an image classification task (Section 7), while the backward GWA seems to be unknown to the best of our knowledge. The backward GWA is not identical to the neural feature ansatz (NFA) but can be seen as an equivariant correction to the NFA (Section 7), and the forward GWA is also unknown to the community. Adding together the forward and backward versions, there are six alignment relations. That these relations simultaneously for any fully connected layer will be referred to as the canonical representation hypothesis (CRH).
@@ 4 Noise-Regularization Balance Leads to Alignment
Now, if additionally either (a) $\mathbb{E}[\Delta W]=0$ (namely, at a local minimum) or (b) $\mathbb{E}[\Delta(WW^{\top})]=0$ holds, the weight will also align with the cross terms between $g_{b}$ and $h_{b}$: $WW^{\top}\propto\mathbb{E}[g_{b}h_{b}^{\top}]+\mathbb{E}[h_{b}g_{b}^{\top}]$, which leads to the effect that the learning term above must also balance with the regularization term. Thus, eventually, the regularization effect will have to be balanced with the gradient noise. Lastly, if both (a) and (b) hold, the alignment between all three matrices emerges: $G_{b}\propto H_{b}\propto Z_{b}$.
@@
One can similarly derive condition for $\mathbb{E}[\Delta(g_{a}g_{a}^{\top})]=0$: $z_{a}(\mathbb{E}[h_{a}g_{a}^{\top}]+\mathbb{E}[g_{a}h_{a}^{\top}])+\eta z_{a}^{2}\mathbb{E}[h_{a}h_{a}^{\top}]=2\gamma\mathbb{E}[g_{a}g_{a}^{\top}]$, where we have defined $z_{a}=\mathbb{E}[\|g_{b}\|^{2}]$. The forward CRH can then be derived if $W^{\top}W$ and $W$ reaches stationarity.
@@
Under Assumption 1, when $\mathbb{E}[\Delta(h_{a}h_{a}^{\top})]=0$, $\mathbb{E}[\Delta(g_{b}g_{b}^{\top})]=0$, $\mathbb{E}[\Delta(WW^{\top})]=0$, and $\mathbb{E}[\Delta(W^{\top}W)]=0$, there exist real-valued constants $c_{1},\ c_{2},\ c_{3},\ c_{4}>0$ such that
@@
A strength of this derivation is that it is oblivious to the loss function, the model architecture, or the type of activation used (as long as the second moments exist). This may explain the wide applicability of the RGA observed in Section 6. The above alignment relations can be seen as a type of fluctuation-dissipation theorem in theoretical physics (Kubo,, 1966), which states that in a driven stochastic dynamics, the fluctuation of the force must balance with the rate of energy loss – a fundamental law first discovered by Einstein, (1905). Prior applications of the fluctuation-dissipation theorem to deep learning have focused on the covariance of the model parameters (Yaida,, 2018; Liu et al.,, 2021) and are not directly relevant to the representations.
@@ 5 CRH Breaking and Polynomial Alignment Hypothesis
Table 1: The reciprocal polynomial relations of the CRH Master Theorem. When one forward relation and one backward relation hold simultaneously, all six matrices are polynomially aligned in a subspace (Theorem 2). Each scaling relationship can be regarded as a possible phase for the layer during actual training. The right panel shows how existing observations about neural networks fit into the phase diagram. A ✓ denotes that this phenomenon is compatible with the specified phase. NC refers to the neural collapse. NFA refers to the neural feature ansatz. CU (correlated update) refers to the (idealization of the) common observation that $h_{a}$ is correlated with $W$ a few steps after training (Everett et al.,, 2024). The llm column shows the compatibility of the scaling relation for transformer observed in Figure 3.
@@
The idea behind the proof of the theorem is that there is a degree of redundancy in the six matrices: every forward relation implies a backward relation and vice versa. As an example, if $Z_{a}\propto H_{a}$, then we also have $Z_{b}^{2}\propto H_{b}$, which can be obtained by multiplying $W$ on the left and $W^{\top}$ on the right.
@@
The last part of the theorem clarifies what it means to perfectly satisfy the CRH: the latent representation is fully compact, where weight and representation are only nonvanishing in the subspaces where the gradient is nonvanishing. Moreover, the weight matrix does nothing but rotates the representation, implying that the information processing is invertible once the CRH is fully satisfied. This is consistent with the observation that once one layer has perfect alignment, all the layers after it also have almost perfect alignment, a sign that the representation cannot be further compressed (Section 6). Therefore, the CRH is consistent with the observation that last layers of neural networks are low-rank and invariant to irrelevant features.
@@
Part (2) of the theorem is especially relevant when the the CRH is only partially satisfied. Depending on which subset of the hypotheses holds well, the learning dynamics and outcome may be classifiable into as many as $642^{6}=$ 64 phases. In different phases, the learning dynamics and the found solution will likely be completely different due to different scaling relations. For example, positive exponents between $Z_{a}$ and $H_{a}$ will imply that the layer is enhancing the principle components of $H_{a}$, while suppressing the lesser features; a negative exponent would imply the converse. Even if we remove the redundancy implied by the theorem, there are still at least $226+{6\choose 2}+1=$ 22 phases. One can also define additional phases according to the ordering of the matching of each relation when the matching is imperfect, which gives $6!=720$ phases, although these ordering phases may not have major influence on training. In many experiments we performed, at least one of the forward relations and one of the backward relations are observed to hold very well (e.g., Figure 8). This means that one is quite likely to observe the power-law relations predicted in Table 1. We also find it common for different layers to be in different phases, even within the same network.
@@
Broadly interpreted, part (2) predicts a power law relation between the spectrum of all six matrices, which is also what we observe in almost all experiments (Section 6). What is quite surprising is that almost all positive exponents we observed are within the range $[1/3,3]$, which is exactly the range of exponents that the theorem predicts (e.g., see Section C.8). Formally, that the $H$, $Z$, and $G$ are polynomially related can be called the “Polynomial Alignment Hypothesis (PAH)” and is a natural extension of the CRH. That scaling relations can be used to characterize different phases of matter is an old idea in science. In physics, phases can be classified according to their scaling exponents, and having a different set of exponents implies that the underlying dynamics and mechanism are entirely different (Pelissetto and Vicari,, 2002). This connection corroborates our physics-inspired proof.
@@
A major remaining question is whether we can find mechanisms such that the CRH breaks. The theory in the previous section implies one primary mechanism that the CRH breaks. For all six alignments to hold, it needs to be the case that both $\mathbb{E}[\Delta W]$, $\mathbb{E}[\Delta Z]$ are zero, but these two conditions may not be easily compatible with each other, as they together imply that $\Delta W$ has zero variance. While this may be possible for some subspaces (e.g., see Ziyin et al., (2024)), it does not hold for every subspace. In fact, it is easy to show that unless the minibatch size is large enough the SGD updates will never reach zero variance that is $\Delta W\neq 0$ even for $t\to\infty$ (see Lemma 8 in Xu et al., 2023b ; see also Xu et al., 2023a ). Thus, the competition between reaching a training loss of zero and the need to reach a stationary fluctuation is a primary cause of the CRH breakage. Because the rank tends to decrease for later layers in networks performing classification, we expect that CRH holds better for later layers than for earlier layers.
This problem is especially troublesome in the layers where $\mathbb{E}[\Delta W]$ has a high rank, which holds true for the earlier layers of the network but not the later layers (Xu et al., 2023a, ). This is consistent with the observation that the CRH holds much better in the latter layers than in the beginning layers. This analysis is consistent with the fact that a stronger alignment is strongly correlated with a more compact representation (Figure 4).
@@ 6 Experiments
We would like to measure how similar and well-aligned the six matrices in the CRH are. Let $A$ and $B$ be two square matrices, each with $d^{2}$ real-valued elements. We use the Pearson correlation to measure the alignment between two matrices (Herdin et al.,, 2005):
@@
Fully Connected Nets (fc1). We also perform a systematic exploration of how hyperparameters of training and model architectures affect the formation of representations. Results presented in Section C.6 show the following phenomena: (1) Gradual Alignment: deeper layers tend to have better alignment; if a layer has an almost perfect alignment ($\alpha\approx 1$), then any layer after it will also have almost perfect alignment (note the similarity of this phenomenon with neural collapse); this is consistent with the tunnel effect discovered in (Masarczyk et al.,, 2024); (b) Critical Depth and Alignment Separation: batch size $B$ affects the alignment significantly; earlier layers have worse alignment as $B$ increases; later layers have better alignment as $B$ increases; wider networks have similar level of alignment as thinner ones for SGD; for Adam, it is more subtle: early layers have worse alignment for a large width, later layers have better alignment for a large width; (c) weight decay $\gamma$ affects the alignment significantly and systematically; larger $\gamma$ always leads to better alignment.

### v2

@@ Abstract
Understanding neural representations will help open the black box of neural networks and advance our scientific understanding of modern AI systems. However, how complex, structured, and transferable representations emerge in modern neural networks has remained a mystery. Building on previous results, we propose the Canonical Representation Hypothesis (CRH), which posits a set of six alignment relations to universally govern the formation of representations in most hidden layers of a neural network. Under the CRH, the latent representations (R), weights (W), and neuron gradients (G) become mutually aligned during training. This alignment implies that neural networks naturally learn compact representations, where neurons and weights are invariant to task-irrelevant transformations. We then show that the breaking of CRH leads to the emergence of reciprocal power-law relations between R, W, and G, which we refer to as the Polynomial Alignment Hypothesis (PAH). We present a minimal-assumption theory proving that the balance between gradient noise and regularization is crucial for the emergence of the canonical representation. The CRH and PAH lead to an exciting possibility of unifying major key deep learning phenomena, including neural collapse and the neural feature ansatz, in a single framework.
@@ 1 Introduction
The success of deep learning is often attributed to its ability to learn meaningful latent representations from data (Bengio et al., 2013). These latent representations, progressively formed as data passes through the network’s layers, are found to encode increasingly abstract features of the input:
@@
In this work, we seek to bridge this gap by introducing the Canonical Representation Hypothesis (CRH). At its core, the CRH posits that neural networks, during training, inherently align their representations with the gradients and weights. The satisfaction and breaking of a subset of CRH equations are found to delineate the universal phases, which are empirically observable scaling relationships between the weights, activations, and gradients. The CRH reveals a striking aspect of representation learning: there may exist a set of universal equations that govern the formation of representations and universal phases which distinguish the layers in modern neural networks, independent of the task, architecture, or loss function. Thus, the CRH provides an useful perspective on how neural networks evolve toward compact and interpretable solutions.
@@
    Proposal and justification of the CRH which states that within a layer, the neuron gradients, latent representation, and parameters are driven into mutual alignment after training, due to noise-regularization balance (Section 3 and 4);
@@ 2 Related Works
Empirical results show that the representations of well-trained neural networks share universal characteristics (Maheswaranathan et al., 2019; Huh et al., 2024; Ziyin et al., 2025). For us, a closely related phenomenon is the neural collapse (NC) (Papyan et al., 2020), which studies how structured and noise-robust low-rank representations emerge in a classification model. The CRH can be seen as a generalization of NC because one can prove that when restricted to certain settings, the CRH is equivalent to the NC (Section 7). Another related phenomenon is the neural feature ansatz (NFA) (Radhakrishnan et al., 2023), which shows that the weight matrices of fully connected layers evolve according to the gradient outer product during training. However, the NFA studies the weight evolution, not the representations. Empirical power-laws are known to exist in large neural networks (Kaplan et al., 2020; Bahri et al., 2024), which relates the model performance to their sizes. The power laws discovered in our work are different because they are reciprocal relations that relate dual objects (R, G, W) to each other rather than to the performance. Other related works are discussed in the context where they become relevant. More related works are discussed in Appendix A.
@@ 3 Canonical Representation Hypothesis
Much recent literature has suggested that the quantities $h_{a},\ h_{b},\ W$ and their gradients are correlated with each other after or throughout training. The neural collapse phenomenon suggests that in a deep overparameterized classifier, $\mathbb{E}[h_{a}h_{a}^{\top}]\propto W^{\top}W$ for the penultimate fully connected layer (Papyan et al., 2020; Xu et al., 2023b; Ji et al., 2021; Kothapalli, 2022; Rangamani et al., 2023). In the study of kernel and feature learning, a primary mechanism of how the neural tangent kernel changes is that after a few steps of update, the representations become correlated with the weights (Everett et al., 2024) and so the quantity $Wh_{a}$ will be significantly away from zero, which also implies a strong relationship between $W^{\top}W$ and $\mathbb{E}[h_{a}h_{a}^{\top}]$. The recent work on neural feature ansatz shows that $W^{\top}W\propto\nabla_{h_{a}}f\nabla_{h_{a}}^{\top}f$ for fully connected networks (Radhakrishnan et al., 2023; Gan & Poggio, 2024). The idea that the latent variables will become correlated with the weight updates is also a central notion in the feature learning literature (Yang & Hu, 2020; Everett et al., 2024).
@@
For the formation of representations in neural networks, the most important relations in Eq. (3)-(5) are perhaps the forward and backward RGA, as they directly relate the representations $\mathbb{E}[hh^{\top}]$ to their gradients with respect to the loss function. Because there is no scientific reason for us to believe that the forward representation is more important than the backward representation or vice versa, one should study both directions carefully. The backward RWA in its general form has not been discussed in the literature but has been implicitly studied in the particular setting of neural collapse, which happens in the penultimate layer of an image classification task (Section 7), while the backward GWA seems to be unknown to the best of our knowledge. The backward GWA is not identical to the neural feature ansatz (NFA) but can be seen as an equivariant correction to the NFA (Section 7), and the forward GWA is also unknown to the community. Adding together the forward and backward versions, there are six alignment relations. That these relations simultaneously hold for any fully connected layer will be referred to as the canonical representation hypothesis (CRH).
@@ 4 Noise-Regularization Balance Leads to Alignment
Now, if additionally either (a) $\mathbb{E}[\Delta W]=0$ (namely, at a local minimum) or (b) $\mathbb{E}[\Delta(WW^{\top})]=0$ holds (Section B), the weight will also align with the cross terms between $g_{b}$ and $h_{b}$: $WW^{\top}\propto\mathbb{E}[g_{b}h_{b}^{\top}]+\mathbb{E}[h_{b}g_{b}^{\top}]$, which leads to the effect that the learning term above must also balance with the regularization term. Thus, eventually, the regularization effect will have to be balanced with the gradient noise. Lastly, if both (a) and (b) hold, the alignment between all three matrices emerges: $G_{b}\propto H_{b}\propto Z_{b}$.
@@
One can similarly derive condition for $\mathbb{E}[\Delta G_{a}]=0$: $z_{a}(\mathbb{E}[h_{a}g_{a}^{\top}]+\mathbb{E}[g_{a}h_{a}^{\top}])+\eta z_{a}^{2}\mathbb{E}[h_{a}h_{a}^{\top}]=2\gamma\mathbb{E}[g_{a}g_{a}^{\top}]$, where we have defined $z_{a}=\mathbb{E}[\|g_{b}\|^{2}]$. The forward CRH can then be derived if $W^{\top}W$ and $W$ reaches stationarity.
@@
Under Assumption 1, when $\mathbb{E}[\Delta H_{a}]=0$, $\mathbb{E}[\Delta G_{b}]=0$, $\mathbb{E}[\Delta(WW^{\top})]=0$, and $\mathbb{E}[\Delta(W^{\top}W)]=0$, there exist real-valued constants $c_{1},\ c_{2},\ c_{3},\ c_{4}>0$ such that
@@
A strength of this derivation is that it is oblivious to the loss function, the model architecture, or the type of activation used (as long as the second moments exist). This may explain the wide applicability of the RGA observed in Section 6. The above alignment relations can be seen as a type of fluctuation-dissipation theorem in theoretical physics (Kubo, 1966), which states that in a driven stochastic dynamics, the fluctuation of the force must balance with the rate of energy loss – a fundamental law first discovered by Einstein (1905). Prior applications of the fluctuation-dissipation theorem to deep learning have focused on the covariance of the model parameters (Yaida, 2018; Liu et al., 2021) and are not directly relevant to the representations.
@@ 5 CRH Breaking and Polynomial Alignment Hypothesis
Table 1: The reciprocal polynomial relations of the CRH Master Theorem. When one forward relation and one backward relation hold simultaneously, all six matrices are polynomially aligned in a subspace (Theorem 2). Each scaling relationship can be regarded as a possible phase for the layer during actual training. The right panel shows how existing observations about neural networks fit into the phase diagram. A ✓ denotes that this phenomenon is compatible with the specified phase. NC refers to the neural collapse. NFA refers to the neural feature ansatz. CU (correlated update) refers to the (idealization of the) common observation that $h_{a}$ is correlated with $W$ a few steps after training (Everett et al., 2024). The llm column shows the compatibility of the scaling relation for transformer observed in Figure 3.
@@
The idea behind the proof is that there is some redundancy in the six matrices: every forward relation implies a backward relation and vice versa. As an example, if $Z_{a}\propto H_{a}$, then we also have $Z_{b}^{2}\propto H_{b}$, which can be obtained by multiplying $W$ on the left and $W^{\top}$ on the right.
@@
Part (4) of the theorem clarifies what it means to satisfy the CRH: the latent representation is fully compact, where weight and representation are only nonvanishing in the subspaces where the gradient is nonvanishing. Moreover, the weight matrix does nothing but rotates the representation, implying that the information processing is invertible once the CRH is fully satisfied. This is consistent with the observation that once a layer has an almost perfect alignment, all the layers after it also have perfect alignment, a sign that the representation cannot be further compressed (Section 6). Therefore, the CRH is consistent with the observation that last layers of neural networks are low-rank and invariant to irrelevant features.
@@
Part (2) is especially relevant when the CRH is broken. Depending on which subset of the hypotheses holds, the learning process may be classifiable into as many as $642^{6}=$ 64 phases. In different phases, the learning dynamics and the found solution will likely be different due to different scaling relations. For example, positive exponents between $Z_{a}$ and $H_{a}$ will imply that the layer is enhancing the principle components of $H_{a}$, while suppressing the lesser features; a negative exponent would imply the converse. Even if we remove the redundancy implied by the theorem, there are still at least $22$ phases. One can also define additional phases according to the ordering of the degree of breaking for each relation, which gives $6!=720$ phases, although these ordering phases may not have a major influence. In many experiments, we performed, at least one of the forward relations and one of the backward relations are observed to hold very well (e.g., see Figure 8). This means that one is quite likely to observe the power-law relations predicted in Table 1. We also find it common for different layers to be in different phases, even within the same network. We discuss more potential meanings and examples in Section D.
@@
Broadly interpreted, part (2) predicts a power law relation between the spectrum of all six matrices, which is also what we observe in almost all experiments (Section 6). What is quite surprising is that almost all positive exponents we observed are within the range $[1/3,3]$, which is exactly the range of exponents that the theorem predicts (e.g., see Section C.8). Formally, that the $H$, $Z$, and $G$ are polynomially related can be called the “Polynomial Alignment Hypothesis (PAH)” and is a natural extension of the CRH. That scaling relations can be used to characterize different phases of matter is an old idea in science. In physics, phases can be classified according to their scaling exponents, and having a different set of exponents implies that the underlying dynamics and mechanism are entirely different (Pelissetto & Vicari, 2002). This connection corroborates our physics-inspired proof.
@@
A major remaining question is whether we can find mechanisms such that the CRH breaks. The theory in the previous section implies one primary mechanism that the CRH breaks. For all six alignments to hold, it needs to be the case that both $\mathbb{E}[\Delta W]$, $\mathbb{E}[\Delta Z]$ are zero, but these two conditions may not be easily compatible with each other, as they together imply that $\Delta W$ has zero variance. While this may be possible for some subspaces (e.g., see Ziyin et al. (2024a)), it does not hold for every subspace. In fact, it is easy to show that unless the minibatch size is large enough the SGD updates will never reach zero variance that is $\Delta W\neq 0$ even for $t\to\infty$ (see Lemma 8 in Xu et al. (2023b); see also Xu et al. (2023a)). Thus, the competition between reaching a training loss of zero and the need to reach a stationary fluctuation is a primary cause of the CRH breakage. Because the rank tends to decrease for later layers in networks performing classification, we expect that CRH holds better for later layers than for earlier layers.
This problem is especially troublesome in the layers where $\mathbb{E}[\Delta W]$ has a high rank, which holds true for the earlier layers of the network but not the later layers (Xu et al., 2023a). This is consistent with the observation that the CRH holds much better in the latter layers than in the beginning layers. This analysis is consistent with the fact that a stronger alignment is strongly correlated with a more compact representation (Figure 4).
@@ 6 Experiments
We would like to measure how similar and well-aligned the six matrices in the CRH are. Let $A$ and $B$ be two square matrices, each with $d^{2}$ real-valued elements. We use the Pearson correlation to measure the alignment between two matrices (Herdin et al., 2005):
@@
Fully Connected Nets (fc1). We also perform a systematic exploration of how hyperparameters of training and model architectures affect the formation of representations. Results presented in Section C.6 show the following phenomena: (1) Gradual Alignment: deeper layers tend to have better alignment; if a layer has an almost perfect alignment ($\alpha\approx 1$), then any layer after it will also have almost perfect alignment (note the similarity of this phenomenon with neural collapse); this is consistent with the tunnel effect discovered in (Masarczyk et al., 2024); (b) Critical Depth and Alignment Separation: batch size $B$ affects the alignment significantly; earlier layers have worse alignment as $B$ increases; later layers have better alignment as $B$ increases; wider networks have similar level of alignment as thinner ones for SGD; for Adam, it is more subtle: early layers have worse alignment for a large width, later layers have better alignment for a large width; (c) weight decay $\gamma$ affects the alignment significantly and systematically; larger $\gamma$ always leads to better alignment.